- (2^2) + 6(5-9)/2 + 48/2 =

I came up with 23.5

Answer by unlockmath(1688)

Hello,
Let's see what I come up with. Hopefully, same as your, right?
- (2^2) + 6(5-9)/2 + 48/2 =
-4-24/2+24=
-4-12+24=
-16+24=8
So the trick is to do the parenthesis first. Keep those negative signs correct also.
RJ Toftness
www.math-unlock.com